Suppose a subdivision on the southwest side of Denver, Colorado, contains 1,800 houses. The subdivision was built in 1983. A sample of 110 houses is selected randomly and evaluated by an appraiser. If the mean appraised value of a house in this subdivision for all houses is $225,000, with a standard deviation of $8,500, what is the probability that the sample average is greater than $226,000?
Answer #1
sample size             n= 110
population std deviation σ= 8500
population size N= 1800
std errror 'σx̅=(σ/√n)*√((N-n)/(N-1))= 785.5075
for normal distribution z score =(X-μ)/σx

probability that the sample average is greater than $226,000 :

probability =P(X>226000)=P(Z>(226000-225000)/785.508)=P(Z>1.27)=1-P(Z<1.27)=1-0.898=0.1020